CRY+VIS Interface Meeting
Date: 2022/6/3 13:00-13:40
Participants: Takahashi, Miyoki, Ikeda, Akutsu, Uchiyama, Hirata, Sato, Node, Ushiba, Yokozawa, Aso
- IX: MN Oplev Laser noise (Preparing for temperature control). Offload.
- EX: Recovery from cooling (Setpoint change, Offload, Health check).
- IY: Leak test. Offload.
- EY: Leak test.
- Type-B: Recovery from FC on BS (OSEM adjustment, Cable touching, 50mHz yaw motion). Visual inspection of BS is necessary. Stand-alone SM driver for SR3. Geophone #2 in SR2.
- Type-Bp: Test of traverser limit SW in PRM. Installation of MSB to BS. Health check of PRs. FC on PR3 and traverser limit SW to PR3 in the late next week.
- Type-C: Alignment of OMC in the next week (Start from Y-arm alignment).
- Replacement of flip cables: EXT and BF temp.
- TM height setting: Following the cooling.
- Oplev: TM Oplev windshield in EX.
- Commissioning in air:
- Commissioning in vacuum: Y-arm lock with Stefan and transmission of X-arm in the next week.
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Tested in IYA. Test in IYV in the next week.
- Heater control: Pending.
- TMS: QPD in EX.
Date: 2022/6/3 14:00-14:25
Participants: Takahashi, Uchiyama, Fabian, Aso, Tanaka, Ushiba, Washimi
- Offloaded the GAS filters in IX, EX, and IY.
- Beam Splitter: after First Contact work last week, the H2 OSEM readout was close to the end of its usable range.
- We found a cable from one of the RM coils touching the security structure. We moved it and clamped it in a slightly different position.
- Then we adjusted the pitch and yaw of the BS using the picomotor and F0 yaw stepper motor respectively.
- We also inspected that nothing around the BF was touching the security structure. Ushiba-san also made a quick test moving up and down the F0 and checking the output of F1 and BF LVDTs.
- Finally, we moved H2 and H3 OSEMs to more suitable positions.
- We measured TFs and corroborated that the system is healthy.
- In SR2, the geophones stopped working, so we checked. The problem was tracked down to the geophone #2 pins powering the preamplfier. There is a short circuit between pins 4 and 5. Next week we will check the in-vacuum cables and, depending on the result, we may want to change the preamplfier or the whole geophone with its pod.
- Health check of PR suspensions.
- Type-A: Y-arm lock.
- Type-B: Geophone in SR2.
- Type-Bp: FC on PR3. Traverser limit SW in PR3.
- Type-C: Alignmenet of OMC.
(Date in Kamioka)
The week from 6/6
- Takahashi: 6/7-10
- Sato: 6/6-10
- Hirata: 6/6-10
- Aso: 6/6-9
The week from 6/13